PLEASE HELP ME!!
Find two numbers whose sum is 25 and the sum of whose reciprocals is 1/4.

Answer by amoresroy(361)

Find two numbers whose sum is 25 and the sum of whose reciprocals is 1/4.
Let x = the 1st number
25-x = the 2nd number
1/x + 1/(25-x)= 1/4
multiply both sides by 4(x)(25-x)
4(25-x) + 4x = x(25-x)
100-4x +4x = 25x -x^2
100 = 25x - x^2
x^2-25x+100 = 0
solve by factoring
(x-20)(x-5) = 0
x-20 = 0
x = 20
x-5 = 0
x = 5
The numbers are 20 & 5